# Evolutionary Generative Optimization: Towards Fully Data-Driven Evolutionary Optimization via Generative Learning

### v1

@@ IV-E Search Dynamics
Finally, to answer our third research question (RQ3) and gain deeper insights into the search dynamics of EvoGO, we conducted a visualization study on a toy problem. Specifically, Fig. 8 investigates the search dynamics of EvoGO by applying it to a two-dimensional Ackley function with a population size of 100. Here, we tracked the transitions from input solutions $\bm{x}_{\text{in}}$ to generated output solutions $\bm{x}_{\text{out}}=\bm{\gamma}(\bm{x}_{\text{in}})$, captured across various generations from the trained generative model $\bm{\gamma}$. Arrows in the visualization represent vectors from $\bm{x}_{\text{in}}$ to $\bm{x}_{\text{out}}$, and the colors of these arrows reflect their norm-2 lengths, as described in the corresponding plot legends. This visual representation clearly demonstrates EvoGO’s ability to precisely estimate both the direction $(\bm{x}_{\text{out}}-\bm{x}_{\text{in}})/\|\bm{x}_{\text{out}}-\bm{x}_{\text{in}}\|$ and the magnitude $\|\bm{x}_{\text{out}}-\bm{x}_{\text{in}}\|$ of steps towards potential optima. This capability not only enhances optimization efficiency but also significantly mitigates mode collapse. Furthermore, $\bm{\gamma}$’s learning progress is illustrated through the grayscale contours of the optimization landscapes captured by the solutions at different generations. Throughout the search process, $\bm{\gamma}$ adaptively hones in on specific regions, showing a refined focus that enhances its search efficacy. This adaptiveness suggests EvoGO’s potential for achieving superlinear convergence given ample data, which is accentuated by the dashed rectangles that delineate the boundaries of solutions generated at different generations.
@@ V Conclusion
This paper presents Evolutionary Generative Optimization (EvoGO), a data-driven evolutionary optimization framework designed to address three critical limitations in existing generative model-based evolutionary algorithms: reliance on handcrafted heuristics, objective misalignment between generative models and evolutionary goals, and underutilization of available data. By eliminating traditional search operators and adopting globally consistent training objectives, EvoGO enables end-to-end learning for optimization with enhanced generality, alignment, and data efficiency. The proposed framework introduces a paired generative architecture that integrates forward and inverse models to produce high-quality candidate solutions, effectively replacing conventional crossover and mutation. A pairwise data construction strategy is employed to enhance training diversity under limited evaluation budgets. Meanwhile, a composite loss function that integrates reconstruction, distributional alignment, and directional guidance ensures efficient model convergence toward promising regions. Extensive experiments across numerical, control, and robotic tasks demonstrate the superior performance of EvoGO over state-of-the-art baselines. It achieves convergence within 10 generations and delivers up to a 134× speedup on complex tasks through GPU-parallelized inference. These findings highlight the potential of fully learnable evolutionary frameworks to generalize across domains without manual design. Future work will focus on enhancing performance under extremely low-data regimes and providing theoretical insights into convergence behavior and optimization dynamics.

### v2

@@ IV-E Search Dynamics
To answer our third research question (RQ3) and gain deeper insights into the search dynamics of EvoGO, we conducted a visualization study on a two-dimensional Ackley function with a population size of 100, as shown in Fig. 7. Specifically, we tracked the transitions from input solutions $\bm{x}_{\text{in}}$ to generated output solutions $\bm{x}_{\text{out}}=\bm{\gamma}(\bm{x}_{\text{in}})$, captured across various generations from the trained generative model $\bm{\gamma}$. Arrows in the visualization represent vectors from $\bm{x}_{\text{in}}$ to $\bm{x}_{\text{out}}$, and their colors reflect their norm-2 lengths, as described in the corresponding plot legends. This visual representation demonstrates EvoGO’s capacity to precisely estimate both the direction $(\bm{x}_{\text{out}}-\bm{x}_{\text{in}})/\|\bm{x}_{\text{out}}-\bm{x}_{\text{in}}\|$ and the magnitude $\|\bm{x}_{\text{out}}-\bm{x}_{\text{in}}\|$ of steps toward potential optima. This capability not only enhances optimization efficiency but also significantly mitigates mode collapse. Furthermore, the learning progress of $\bm{\gamma}$ is illustrated through the grayscale contours of the optimization landscapes captured by the solutions at different generations. Throughout the search process, $\bm{\gamma}$ adaptively hones in on specific regions, thereby demonstrating a refined focus that enhances its search efficacy. This adaptiveness suggests EvoGO’s potential for achieving superlinear convergence given ample data, which is accentuated by the dashed rectangles that delineate the boundaries of solutions generated at different generations.
@@ V Conclusion
This paper presented Evolutionary Generative Optimization (EvoGO), a data-driven evolutionary optimization framework. This framework was specifically designed to address three critical limitations inherent in existing generative model-based evolutionary algorithms: reliance on handcrafted heuristics, objective misalignment between generative models and evolutionary goals, and underutilization of available data. By eliminating traditional search operators and adopting globally consistent training objectives, EvoGO enables robust end-to-end learning for optimization, offering enhanced generality, objective alignment, and data efficiency.
The proposed framework is underpinned by a paired generative architecture that integrates forward and inverse models to produce high-fidelity candidate solutions, thereby effectively replacing conventional crossover and mutation operators. To enhance training diversity under limited evaluation budgets, we employed a pairwise data construction strategy. Furthermore, a composite loss function integrates reconstruction, distributional alignment, and directional guidance to ensure efficient model convergence toward promising objective regions.
Extensive experiments conducted across numerical, control, and robotic tasks demonstrated the superior performance of EvoGO when compared against state-of-the-art baselines. Specifically, EvoGO consistently achieved convergence within 10 generations. This performance, combined with GPU-parallelized inference, delivered up to a $134\times$ speedup on computationally complex optimization tasks.
These results underscore the substantial potential of fully data-driven evolutionary frameworks to generalize effectively across diverse domains without requiring manual heuristic design. Future work will initially focus on enhancing robustness under extremely low data regimes and providing theoretical insights into the framework’s convergence behavior and optimization dynamics. In addition, future extensions will explore offline pre-training and cross-task transfer to further improve data efficiency in settings where online evaluation is limited.